# Accelerated Real-time Cine and Flow Under In-magnet Staged Exercise

### v2

@@ 3 Results
3.1 Results from healthy subjects
@@
In Figure 2A, the cardiac function parameters extracted from the short-axis RT cine images are displayed, tracing the progression from rest to three stages of exercise for each volunteer. In summary, the EDV remained relatively consistent from rest through the exercise stages in both ventricles. In contrast, the ESV decreased as exercise intensity and HR increased, indicating more vigorous cardiac contractions during exercise stress. Consequently, the SV typically increased under stress, which is consistent with the literature [17]. However, the pattern of increase in SV varied among volunteers, ranging from a rapid initial rise to plateauing in some, and a more gradual increase in others. The consistent elevation in CO was attributable to the increased SV and HR. Both HR and APHR trends extracted from the cine images (not shown) were similar to the ones extracted from the flow images.
@@
In Figure 2B, the aortic and pulmonic flow parameters, tracked from rest through three stages of exercise for each volunteer, are depicted. The NFF values demonstrate an overall increase with exercise, although this rise is less pronounced compared to the biventricular stroke volumes. However, CO exhibits a similar trend of increasing with exercise intensity. The peak velocities for both arteries also show an upward trend with exercise. Across all volunteers, the maximum heart rates (recorded from cine or flow) ranged from 92 to 152 bpm and the maximum APHR ranged from 46% to 81%. On average, the subjects were able to achieve a maximum APHR value of 70%. The perceived exertion, measured using BRPE, ranged from 6 to 16, with an average value of 9.3. For the subjects who quit before reaching the $60$ W workload, the most commonly expressed reasons were shortness of breath, leg fatigue, and overall exhaustion.
@@
Representative cine and flow images at rest and various stages of exercise of a healthy volunteer are shown in Figure 3. The cine images show a clear delineation of the blood-myocardium boundary, with some marginal blurring at $60$ W due to the residual motion. The end-systolic frames exhibit stronger contractility, especially at $60$ W. The flow images are mostly free of motion artifacts, with the $60$ W magnitude images showing marginal blurring. Note that a higher value of velocity encoding gradient (VENC) was used during exercise; therefore, the velocities are scaled differently for images collected during exercise compared to the resting images.
@@
Figure 4 displays the repeat assessment of biventricular functional parameters from RT cine images at rest and the workload of $40$ W. Similarly, Figure 5 presents the repeat assessment of hemodynamic parameters from RT flow images at the same conditions. Overall, both AP and EE beats demonstrate a high level of repeatability. However, in terms of the concordance correlation coefficient (CCC) and normalized mean absolute error (NMAE), EE beats show a marginal but consistent improvement over AP beats; see Figure 6. NMAE, inferred from the two repeats, $x_{1}$ and $x_{2}$, is calculated as $\mathrm{NMSE}=100\times(|x_{1}-x_{2}|)/(\frac{1}{2}|x_{1}+x_{2}|)$.
3.2 Results from patients
@@
The RT cine and flow data were reconstructed from six CMR patients. In all cases, the image quality, after coil reweighting, was deemed adequate to perform cardiac function and hemodynamic quantification. Figure 7 shows biventricular EF and CO values at rest, during exercise, and post-exercise recovery. Due to the differences in the workload increments and widely different exercise capacities, it was not possible to average the data across patients. Except for one patient, where inadequate spatial coverage due to a user error led to LV-RV mismatch, there was a reasonable agreement between CO and EF from LV and RV. The perceived exertion ranged from $6$ to $19$ on the BRPE scale, with an average value of $9.3$. Although less comprehensive compared to the healthy volunteer study, these results indicate the feasibility of implementing ExCMR in a clinical setting.
@@
Figure 8 shows a representative frame from a patient fatigued at the workload of $20$ W. Both RT cine images show clear delineation of blood and myocardium during the systolic and diastolic phases. The flow images are also mostly free of artifacts, with the $20$ W magnitude image showing slight blurring. The phase images, however, clearly capture the flow in the ascending and descending aorta. Note that a higher value of VENC was used at $20$ W; therefore, the velocities are scaled differently in the $20$ W image compared to the resting image.
The artifact suppression method based on coil reweighting was applied to all RT cine data. In cases, where cine images had little to no artifact, coil reweighting had minimal impact on the image quality or quantification. However, in cases where image artifacts were severe, image interpretation or quantitative analysis was not feasible without coil reweighting. Figure 9 shows one such example, where images without coil reweighting exhibit severe artifacts, which are mostly mitigated after coil reweighting. As highlighted in Table 3, the reweighting procedure had no significant ($p=0.4$) impact on the image quality collected at rest but offered a significant improvement ($p<0.01$) for the images collected during exercise. Also, all images, after coil reweighting, received a score of 3 or higher.
Rest BC Rest AC Exercise BC Exercise AC
Average Score 4.78 4.85 3.18* 4.40*
Worst Score 3 4 1 3
Table 3: Subjective scoring of RT cine both before (BC) and after (AC) coil reweighting. The average was computed over 23 image series and two readers. *The pairwise difference between BC and AC was statistically significant ($p<0.01$) for images collected during exercise.
@@ 4 Discussion
Advancements in ExCMR technology are pivotal for enhancing the diagnosis and management of cardiovascular diseases. Specifically, ExCMR with in-magnet exercise allows dynamic monitoring of quantitative cardiac function during exercise, revealing functional impairments that are not apparent during resting cardiovascular exams. Because breath-holding is generally not feasible during exercise, free-breathing RT imaging is often used to collect data. Technical challenges associated with ExCMR with in-magnet exercise include image quality degradation due to motion artifacts and limited spatial and temporal resolution. In this work, we demonstrate the feasibility of in-magnet ExCMR for measuring biventricular cardiac function and hemodynamics during staged exercise in both healthy subjects and patients.
@@
To ensure adequate spatial and temporal resolutions, we employed high acceleration rates. This was facilitated by variable density Cartesian sampling and parameter-free CS-based inline reconstruction. To further enhance image quality, we applied our recently developed coil reweighting technique that effectively minimizes motion artifacts caused by the physical movement of the MRI receive coils. Additionally, we leveraged our newly proposed technique to extract cardiac and respiratory signals directly from the imaging data, eliminating the need for ECG or bellows. This technique allows for consistent selection of end-expiratory heartbeats, thereby minimizing the impact of through-plane motion and improving the repeatability of imaging parameters. In summary, we are able to address the aforementioned limitations for in-magnet ExCMR by (i) achieving high acceleration rates of $8-9$ in RT cine and $16-19$ in RT flow, leading to the spatial resolution of $1.7-2.4$ mm for cine and $2-2.5$ mm for flow, (ii) reducing scan time such that a full short-axis stack can be acquired in less than two minutes, (iii) providing whole heart coverage, (iv) developing data processing that is independent of ECG, (v) providing biventricular quantification, (vi) yielding hemodynamic assessment, (vii) developing inline image reconstruction, (viii) establishing scan-rescan repeatability, (ix) suppressing motion artifacts by coil reweighting, (x) enabling images across multiple exercise intensities, and (xi) mitigating sensitivity to through-plane motion by using end-expiratory beats for quantification.
With the ability to quantify volumes and function at each stage, we showed that the physiologic changes in cardiac function and hemodynamic parameters from rest through various stages of exercise in healthy individuals are predominately through decreasing in ESV and increasing in HR. The values of ESV, both for LV and RV, consistently decreased with increased contractility under stress. The values of EDV remain relatively stable across different exercise intensities. As a result, both SV and EF gradually increase with exercise. The increase in CO is driven largely by the increase in HR. This biventricular increase is observed in both cine and flow images.
For scan-rescan repeatability, we showed that selecting an EE beat, compared to an AP beat, results in a modest but consistent improvement in repeatability. With the sole exception of MPA CO, NMAE from all parameters from the EE beat is less than 10% (Figure 6). In contrast, the NMAE exceeds the 10% threshold four times for AP. A similar trend is observed in CCC. Relatively poor CCC and NMAE values for MPA CO can be attributed to HR variations between the two repeats and the small sample size.
@@
Finally, we are able to show the impact of coil reweighting on image quality. Without applying coil reweighting, some of the cine images are not interpretable or suitable for quantification. With coil reweighting, most of the artifacts are eliminated. Note, the coil reweighting comes at the cost of suppressing contribution from certain coil elements, which can lead to partial loss of signal in some areas of the image. However, we did not observe a case where the benefit of suppressing artifacts was outweighed by the loss of intensity. As highlighted by scores in Table 3, coil reweighting improved images collected during exercise, without negatively impacting images collected at rest.
@@
The limitations of this study include a small sample size. This particular protocol had an extended exercise duration, primarily driven by the inclusion of 4D flow acquisition at each exercise stage. If limited to RT cine and flow measurements, it is feasible to complete scanning at each stage within three minutes and dramatically shorten the exercise protocol to 15 minutes for a 5-staged exercise protocol similar to the Bruce protocol [33]. The additional limitations include the highest achieved heart rate in the volunteers and patients was 152 bpm and 123 bmp, respectively. Future research directions will include further optimizing the ExCMR protocol and improving temporal resolution to be able to image at maximal exercise.
@@ 5 Conclusion
In this preliminary study, we demonstrated the feasibility of ExCMR with in-magnet exercise for quantitative cardiac function evaluation in both healthy volunteers and patients, under multi-stage exercise. The highly accelerated RT imaging was facilitated by a parameter-free CS, overcoming the traditional challenges of limited resolution. Additional key innovative aspects of our technique are the incorporation of a coil reweighting technique and the selection of end-expiratory heartbeats, which resulted in significantly improved image quality and enhanced repeatability of cardiac function measurements.
@@
Refer to caption ABCD
Figure 9: Cine images from one of the patients without coil reweighting (A and C) and with coil reweighting (B and D). Strong motion artifacts are observed in A and C, which are effectively removed in B and D. The images at the bottom represent $\mathrm{x}$-$\mathrm{t}$ profile along the dashed red line passing through the heart. The data were collected during exercise at $40$ W.
Supporting Movie S1: The movie shows a 6 s cine image series from a healthy volunteer collected under $60$ W of exercise. The curve at the bottom represents the respiratory motion, with the moving cross representing the current frame.
@@
  • Ethics approval and consent: For the human subject data, approval was granted by the Institutional Review Board (IRB) at The Ohio State University (2019H0076). Informed consent to participate in the study and to publish results was obtained from all individual participants.

### v3

@@ 3 Results
3.1 Quantification results from healthy subjects
@@
The image quality, after coil reweighting, was deemed adequate to perform quantification in all cases. In Figure 3A, the cardiac function parameters extracted from the short-axis RT cine images are displayed, tracing the progression from rest to three stages of exercise for each volunteer. In summary, the EDV remained relatively consistent from rest through the exercise stages in both ventricles. In contrast, the ESV decreased as exercise intensity and HR increased, indicating more vigorous cardiac contraction during exercise. Consequently, the SV typically increased under stress, which is consistent with the literature [17]. However, the pattern of increase in SV varied among volunteers, ranging from a rapid initial rise to plateauing in some, and a more gradual increase in others. The consistent elevation in CO was attributable to the increased SV and HR. The HR and APHR trends extracted from the cine images (not shown) were similar to the ones extracted from the flow images discussed below.
@@
In Figure 3B, the aortic and pulmonic flow parameters, tracked from rest through three stages of exercise for each volunteer, are depicted. The NFF values demonstrate an overall increase with exercise, although this rise is less pronounced compared to the biventricular stroke volumes. Nonetheless, CO exhibits a similar trend of increasing with exercise intensity. The peak velocities for both arteries also show an upward trend with exercise. Across all volunteers, the maximum heart rates (recorded from cine or flow) ranged from 92 to 152 beats per minute (bpm) and the maximum APHR ranged from 46% to 81%. On average, the subjects were able to achieve a maximum APHR value of 70%. The perceived exertion, measured using BRPE, ranged from 9 to 16, with an average value of 12.3. For the subjects who quit before reaching the $60$ W workload, the most commonly expressed reasons were shortness of breath, leg fatigue, and overall exhaustion.
The Bland-Altman analysis demonstrating the consistency of LV and RV stroke volumes inferred from cine and flow imaging is shown in Figure 4. This analysis includes data pooled from rest and all exercise stages. Cine-based LV vs. RV SV exhibits a small bias (4.2 mL/m2) and a low mean absolute difference (MAD) (5.3 mL/m2), while aortic vs. pulmonic NFF shows an even lower bias (0.2 mL/m2) and MAD (4.6 mL/m2) values. Comparisons between cine- and flow-derived volumes show slightly higher discrepancies, with LV SV vs. AAo NFF yielding a bias of 4.9 mL/m2 and MAD of 7.9 mL/m2, and RV SV vs. MPA NFF showing a bias of 0.9 mL/m2 and a MAD of 5.9 mL/m2. The weaker agreement between cine and flow may be attributed to: (i) differences in imaging parameters, (ii) mild mitral and tricuspid regurgitation commonly observed in healthy subjects [33], (iii) coronary flow, (iv) inaccuracies in background phase correction in flow imaging, (v) inclusion of trabeculation in cine segmentation [34], which may represent a larger part of ESV during exercise, and (vi) heart rate variations between cine and flow. Nonetheless, the bias and MAD remain below 5 mL/m2 and 8 mL/m2, respectively, across all four comparisons, indicating overall consistency.
@@
Representative cine and flow images at rest and various stages of exercise from a healthy volunteer are shown in Figure 5. The cine images show a clear delineation of the blood-myocardium boundary, with some marginal blurring at $60$ W due to the residual motion. The end-systolic frames exhibit stronger contractility, especially at $60$ W. The flow images are mostly free of motion artifacts, with the $60$ W magnitude images showing marginal blurring. Note that a higher value of velocity encoding gradient (VENC) was used during exercise; therefore, the velocities are scaled differently for images collected during exercise compared to the resting images.
@@
Figure 6 displays the repeat assessment of biventricular functional parameters from RT cine images at rest and the workload of $40$ W. Similarly, Figure 7 presents the repeat assessment of hemodynamic parameters from RT flow images at the same conditions. Overall, both AP and EE beats demonstrate a high level of repeatability. However, as shown in Figure 8, in terms of CCC and NMAE, EE beats show a marginal but consistent improvement over AP beats. With the sole exception of MPA CO, NMAE from all parameters from the EE beat is less than 10%. In contrast, the NMAE exceeds the 10% threshold four times for AP. A similar trend is observed in CCC.
3.2 Quantification results from patients
@@
The RT cine data were reconstructed from six patients. In all cases, the image quality, after coil reweighting, was deemed adequate for performing cardiac function. Figure 9 shows biventricular EF and CO values at rest, during exercise, and post-exercise recovery. Due to the differences in the workload increments and widely different exercise capacities, it was not possible to average the data across patients. Except for one patient, where inadequate spatial coverage due to a user error led to LV-RV mismatch, there was a reasonable agreement between CO and EF from LV and RV. Across all patients, the maximum heart rates (recorded from cine) ranged from 91 to 123 bpm and the maximum APHR ranged from 62% to 81%. On average, the patients were able to achieve a maximum APHR value of 69%. The perceived exertion ranged from $13$ to $19$ on the BRPE scale, with an average value of $16.2$.
@@
Figure 10 shows a representative frame from a patient fatigued at the workload of $20$ W. Both RT cine images show clear delineation of blood and myocardium during the systolic and diastolic phases. The flow images are also mostly free of artifacts, with the $20$ W magnitude image showing slight blurring. The phase images, however, clearly capture the flow in the ascending and descending aorta. Note that a higher value of VENC was used at $20$ W; therefore, the velocities are scaled differently in the $20$ W image compared to the resting image. Although less comprehensive compared to the healthy volunteer study, these results indicate the feasibility of implementing ExCMR in a clinical setting.
3.3 Impact of coil reweighting
The artifacts scores are reported in Table 3. The difference was not significant at rest ($p=0.14$) but was significant ($p<0.01$) under exercise. The inter-reader agreement, as assessed using the weighted Kappa statistic, was 0.653, indicating substantial agreement between the two readers. Figure 11 shows examples from a healthy subject and a patient, where images without coil reweighting exhibit severe artifacts; these artifacts are mostly mitigated after coil reweighting.
Rest CR-- Rest CR$+$ Stress CR-- Stress CR$+$
Artifacts ($n=16$) 4.81 4.91 3.61 4.58
Table 3: Artifact levels from ten healthy subjects and six patients before (CR--) and after (CR$+$) coil reweighting. Each reported number represents an average over subjects, two views (short-axis and two-chamber), and two readers.
@@ 4 Discussion
ExCMR with in-magnet exercise allows dynamic monitoring of cardiac function during exercise, revealing functional impairments that are not apparent during resting cardiovascular exams. Because breath-holding is generally not feasible during exercise, free-breathing RT imaging is often used to collect data. Technical challenges associated with imaging during exercise include image quality degradation due to motion artifacts and limited spatial and temporal resolution. In this work, we demonstrate the feasibility of in-magnet ExCMR for measuring biventricular cardiac function and hemodynamics during staged exercise in both healthy subjects and patients.
@@
To ensure adequate spatial and temporal resolutions, we employed high acceleration rates. This was facilitated by variable density Cartesian sampling and parameter-free CS-based inline reconstruction. To further enhance image quality, we proposed and applied a coil reweighting technique that effectively minimizes motion artifacts caused by the physical movement of the MRI receive coils. Additionally, we leveraged our newly proposed technique to extract respiratory signal directly from the imaging data. This technique allows for consistent selection of end-expiratory heartbeats, thereby minimizing the impact of through-plane motion and improving the repeatability of imaging parameters. In summary, we are able to address the aforementioned limitations for in-magnet ExCMR by (i) achieving high acceleration rates of $8-9$ in RT cine and $16-19$ in RT flow, leading to the spatial resolution of $1.7-2.4$ mm for cine and $2-2.5$ mm for flow, (ii) reducing scan time such that a full short-axis stack can be acquired in less than two minutes, (iii) providing whole heart coverage, (iv) developing data processing that is independent of ECG, (v) providing biventricular quantification, (vi) yielding hemodynamic assessment, (vii) developing inline image reconstruction, (viii) establishing scan-rescan repeatability, (ix) suppressing motion artifacts by coil reweighting, (x) enabling images across multiple exercise intensities, and (xi) mitigating sensitivity to through-plane motion by using end-expiratory beats for quantification.
@@
A key component of our framework, coil reweighting, suppressed motion artifacts and enabled quantification from highly accelerated RT imaging across varying exercise intensities. Without coil reweighting, some of the cine images were not interpretable or suitable for quantification. With coil reweighting, most of the artifacts were eliminated. Equally importantly, coil reweighting did not negatively impact images collected at rest. Compared to coil rejection, which can result in a more significant signal loss in certain image areas, coil reweighting offers a softer approach. Nevertheless, coil reweighting suppresses contributions from certain coil elements, which can lead to partial loss of image intensity. However, we did not observe a case where the benefit of suppressing artifacts was outweighed by the loss of intensity.
The performance of the proposed coil reweighting approach depends on the user-defined values of $K$ and $M$. The parameter $K$ represents the number of coil elements that do not experience significant movement. In our MRI experiments, there were $K_{\sf b}=18$ body coil elements and $K_{\sf s}=12$ supine coil elements. Since the supine coil elements are fixed in position and only a fraction of body coil elements have significant movement, we found that setting $K$ equal to or slightly larger than $K_{\sf s}$ is a reasonable choice. In our experience, selecting a value in the range $K_{\sf s}\leq K\leq 20$ gives comparable results. The other parameter $M$ can be considered as an upper limit on the number of coil elements that experience significant movement. When we tracked the change in $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ from clean images at rest to degraded images under exercise, we observed that typically 2 to 8 coil elements had a marked increase in $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$. We conservatively selected $M=10$ based on this observation. In our experience, $8\leq M\leq K_{\sf b}$ gives comparable results.
@@
The limitations of this work include a small sample size, primarily comprising healthy subjects with low average BMI. All studies were conducted on a 3T scanner, and extending this work to lower field strengths may result in reduced image quality. The bore size and length of the magnet also pose challenges; some larger patients may not be able to exercise comfortably inside the magnet. Additionally, this particular protocol had an extended exercise duration, primarily driven by the inclusion of 4D flow acquisition at each exercise stage. If limited to RT cine and flow measurements, it is feasible to complete scanning at each stage within three minutes and dramatically shorten the exercise protocol to 15 minutes for a 5-staged exercise protocol similar to the Bruce protocol [35]. Furthermore, the impact of coil reweighting on cardiac function quantification for imaging at rest, where a breath-held reference is available, remains to be validated. Moreover, although designed for a single set of sensitivity maps, our coil-reweighting method could be extended to multiple sets by incorporating set-specific maps into Equation (1) and summing over them. However, its effectiveness with multiple sets of sensitivity maps has yet to be validated. Another limitation is that the highest achieved heart rate in the volunteers and patients was 152 bpm and 123 bpm, respectively. Future research directions include further optimizing the ExCMR protocol to achieve higher levels of exertion, improving temporal resolution to enable RT imaging at maximal exercise, and extending coil reweighting to 3D applications.
@@ 5 Conclusion
In this preliminary study, we demonstrated the feasibility of ExCMR with in-magnet exercise for quantitative cardiac function evaluation in both healthy volunteers and patients under multi-stage exercise. The highly accelerated RT imaging was facilitated by a parameter-free CS, overcoming the traditional challenges of limited resolution. Additional key innovative aspects of our technique are the incorporation of coil reweighting and the selection of end-expiratory heartbeats, which resulted in significantly improved image quality and enhanced repeatability of cardiac function measurements.
@@
Figure 2: (A) Pipeline to generate time-averaged coil images $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ and their re-estimated versions, $\{\tilde{\boldsymbol{x}}_{i}\}_{i=1}^{N}$. (B) $\boldsymbol{x}_{i}$, $\tilde{\boldsymbol{x}}_{i}$, and their pixel-wise difference from two coils, one with significant motion artifacts (Coil-4) and one without (Coil-11), are shown. The horizontal lines in Coil-11 are likely from a faint RF interference. This dataset was collected during in-magnet exercise. (C) The residual signal $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ for the $i^{\text{th}}$ coil, with the values from Coil-4 and Coil-11 highlighted with red and green arrows, respectively. For reference, $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ values of all coils were below $6\times 10^{-4}$ at rest (not displayed).
@@
Figure 4: Bland-Altman plots demonstrating the consistency of LV and RV stroke volumes from cine and flow imaging. This analysis includes data pooled from rest and all exercise stages. Top-left: LV and RV SV from cine. Top-right: AAo NFF and MPA NFF from flow imaging. Bottom-left: LV SV from cine and AAo NFF from flow imaging. Bottom-right: RV SV from cine and MPA NFF from flow imaging.
@@
Figure 11: Exercise images reconstructed without (CR--) and with (CR$+$) coil reweighting. Temporal profiles along the green dashed lines are shown, with green arrows highlighting the motion artifacts. The top row displays two views from a healthy volunteer, and the bottom row shows two views from a patient.
Supporting Movie S1: The movie shows a 6 s cine image series from a healthy volunteer collected under $60$ W of exercise. The curve at the bottom represents the respiratory motion, with the moving cross highlighting the current frame.
@@
  • Ethics approval and consent: For the human subject data, approval was granted by the Institutional Review Board (IRB) at The Ohio State University (2019H0076). Informed consent to participate in the study and publish results was obtained from all individual participants.